What is 4x + 2(2x-5) = 8?

Answer by fractalier(6550)

Okay from
4x + 2(2x-5) = 8
we get
4x + 4x - 10 = 8
8x - 10 = 8
8x = 18
x = 18/8 = 9/4 = 2.25